OpenStudy (anonymous):

Find the derivative of the function. G(y) = (y − 2)^4/(y^2 + 4y)^9

OpenStudy (anonymous):

\[G(y)=\frac{(y-2)^4}{(y^2+4y)^9}=\frac{f(x)}{g(x)^9}=\frac{f(x)}{h(x)}\] So we will need the chain and quotient rule. We know that: \[\frac{d}{dx}(f\circ g)(x)=f'(g(x))g'(x)\phantom{...}(chain\phantom{.}rule)\] We also know that: \[\frac{d}{dx}\frac{f(x)}{g(x)}=\frac{f'(x)g(x)-g'(x)f(x)}{[g(x)]^2}\] You remember these?

OpenStudy (anonymous):

which one would I use first?

OpenStudy (anonymous):

Well we have initially: \[\frac{d}{dy}G(y)=\frac{d}{dx}\frac{f(y)}{h(y)}=\frac{f'(y)h(y)-h'(y)f(y)}{h(y)^2}\] Then we can make the substitution: \(h(y)=[g(y)]^9\) \[\frac{d}{dx}G(y)=\frac{f'(y)[g(y)]^9-[g(y)^9]'f(y)}{[g(y)^9]^2}=\frac{f'(y)[g(y)]^9-[g(y)^9]'f(y)}{g(y)^{18}}\] Now, it's no issue subbing \(g(y)\) into \(f'(y)[g(y)^9]\) but before we do that into the second term. We can simplify \([g(y)^9]'\) into \(9g(y)^8g'(y)\) So then i'll simplify a little bit: \[\frac{f'(y)[g(y)]^9-[g(y)^9]'f(y)}{g(y)^{18}}=\frac{f'(y)(y^2+4y)^9-9(y^2+4y)^8g'(y)(y-2)^4}{(y^2+4y)^{18}}\]

OpenStudy (anonymous):

Now all that's left to do is to calculate f'(y) and g'(y) and we're done! Following so far?

OpenStudy (anonymous):

Yes, then after that it the chain rule will be used?

OpenStudy (anonymous):

Ehh well no! There's no chaining needed to find f'(y) or g'(y)! Where I used the chain rule was in deriving \(g(y)^9\) and obtaining \(\frac{d}{dx}g(y)^9=9g(y)^8g'(y)\)

OpenStudy (anonymous):

So we have: \[f(y)=(y-2)^4\rightarrow f'(y)=4(y-2)^3\] And \[g(y)=y^2+4y\rightarrow g'(y)=2y+4\] So we can substitute both these into the big equation. Therefore: \[\frac{d}{dx}G(y)=\frac{4(y-2)^3(y^2+4y)^9-9(y^2+4y)^8(2y+4)(y-2)^4}{(y^2+4y)^{18}}\] And we can simplify a bit: \[=\frac{4(y-2)^3}{(y^2+4y)^9}-\frac{18(y+2)(y-2)^4}{(y^2+4y)^{10}}\]

OpenStudy (anonymous):

And I mean if you REALLY wanted to go ahead, then you would simplify it to this: \[\eqalign{\frac{d}{dy}G(y)&=(y-2)^3\left[\frac{4}{(y^2+4y)^9}-\frac{18(y+2)(y-2)}{(y^2+4y)^{10}}\right] \\ &=(y-2)^3\left[\frac{4}{(y^2+4y)^9}-\frac{18(y^2-4)}{(y^2+4y)^{10}}\right] }\]

OpenStudy (anonymous):

It's a little nicer :)

OpenStudy (anonymous):

AAAND THAT was a messy question haha

OpenStudy (anonymous):

Engineering Math for ya!

OpenStudy (anonymous):

Haha just you wait till multiple integrals and surface area and volume of 3d figures using integration...THATS MESSY lol 's all part of a good experience